# Learning and Transferring Physical Models through Derivatives

### v2

@@ Appendix D Additional Results and Experiments for Section 4
D.1 Allen-Cahn equation
@@ D.2 Continuity equation
The original solution was calculated using finite volumes (Ferziger & Peric, 2001) on a grid with $\Delta x=\Delta y=0.01$ in the domain [1.5,1.5][-1.5,1.5] and time discretization of $\Delta t=0.001$ for $t\in[0,10]$. For the points used during training, we downsampled the full grid by a factor of $10$ for each spatiotemporal coordinate. In the testing phase, we test the model on the full grid, without downsampling. This way, we can test which model can best learn the solution on a finer grid, while being consistent with the equation. The empirical derivatives are calculated with $h=0.01$ on each component.
@@
D.3 Navier-Stokes: Kovasznay flow
@@ D.4 Pendulum
D.4.1 Differentiability with respect to Initial Conditions
@@ D.4.2 HNN and LNN setup
D.4.3 Results with empirical derivatives
@@ D.4.4 Conservative pendulum
D.5 Parametric Allen-Cahn Equation
@@
D.5.1 Experiment with One parameter
Setup.
The MLPs have $4$ hidden layers with $50$ units each. The models take as input the coordinate and parameter of the equation $(x,y,\xi)$ and output the solution for the Allen-Cahn equation with parameter $\xi$, that is $\hat{u}(x,y;\xi)$. The networks are trained with batch size $100$ with the Adam optimizer for $50$ epochs, then are further optimized for another $50$ epochs with the BFGS optimizer on the whole dataset, similarly to Penwarden et al. (2023). The solution on $1000$ random collocation points is calculated for $16$ values of $\xi$. A train/validation/test split is applied, with $8$ parameters for training, $2$ for validation, and $6$ for testing. Testing values of $\xi$ are chosen to explore regions of the parameter space not seen during training and validation. The training and validation datasets are used for tuning the hyperparameters of the models, that is the learning rate and the weights of the loss terms in equation 16.
@@ Results.
D.5.2 Experiment with Two Parameters
@@
This Section provides additional results and information on the setup for the double parametric Allen-Cahn experiment in Section 4.2. The MLPs have $4$ hidden layers with $50$ units each. The models take as input the coordinate and parameters of the equation $(x,y,\xi_{1},\xi_{2})$ and output the estimate of $u$ with parameters $(\xi_{1},\xi_{2})$, that is $\hat{u}(x,y;\xi_{1},\xi_{2})$. The networks are trained with batch size $100$ with the Adam optimizer for $100$ epochs, then are further optimized for another $100$ epochs with the BFGS optimizer on the whole dataset, similarly to Penwarden et al. (2023). The solution on $1000$ random collocation points is calculated for $25$ couples of $(\xi_{1},\xi_{2})$. A train/validation/test split is applied, with $15$ couples of parameters for training, $5$ for validation, and $5$ for testing. The training and validation datasets are used for tuning the hyperparameters of the models, that is the learning rate and the weights of the loss terms in equation 16.
@@
Appendix E Additional Results and Experiments for Section 5
E.1 Higher Order Derivatives
@@
We consider the Korteweg-de Vries (KdV) equation, a third-order non-linear PDE (equation (KdV) in table 8), with $\nu=0.0025$. The IC is $u(0,x)=\cos(\pi x)$ with periodic BC for $u$ and $u_{x}$. The reference solution is first obtained using the Scipy (Virtanen et al., 2020) solver with Fast Fourier Transforms on a grid with $\Delta x=\Delta t=0.005$. Then, a PINN is carefully trained on equation (KdV) in table 8 and is treated as the teacher model to distill knowledge from. For the teacher model, we used an MLP with $9$ layers of $50$ units, $\tanh$ activation, trained for $200$ epochs with batch size $64$ and Adam optimizer. After the teacher’s training, we save a dataset of its outputs $\hat{u}_{T}(x,t)$, its derivatives $\mathrm{D}\hat{u}_{T}(x,t)$, and its hessian matrix $\mathrm{D}^{2}\hat{u}_{T}(x,t)$ on the training collocation points. As always, partial derivatives are calculated via AD Baydin et al. (2018). Then, the student networks are trained with the specific distillation loss of the method, e.g. $\mathrm{D}\hat{u}_{S}-\mathrm{D}\hat{u}_{T}$ for DERL, and the original BC and IC. Since this is a third-order PDE, we are also interested in understanding the impact of higher-order derivatives on the performance, both in terms of the PDE residual and in terms of matching the true solution $u$ or the BC. We implemented Hessian learning (HESL), which learns the Hessian matrices $\mathrm{D}^{2}\hat{u}_{T}$ of the teacher model. We also explored its combination with DERL (DER+HESL, which learns both $\mathrm{D}\hat{u}_{T}$ and $\mathrm{D}^{2}\hat{u}_{T}(x,t)$) and Sobolev (SOB+HESL). In this last case, we use the approximation of the Hessian matrix as suggested in Czarnecki et al. (2017), while for HESL we use the full Hessian of the network, at the cost of an increase in computational time (see Section A.1 for further details). Each methodology has been tuned individually to find the hyperparameters that provide the best approximation of the true teacher’s solution $\hat{u}_{T}$.
@@ Results.
E.2 Neural Conservation Laws
@@
E.3 Transfer to a Larger Timespan
@@
    A second model is trained for $30.000$ steps and the Adam optimizer with the same loss as in equation 7 that consists of (a) a PINN loss term to find the solution to the equation in the second interval $0.35,0.7]\times[-1,1]$, (b) a distillation loss on the dataset from step 1, that is on points of the region $0,0.35]\times[-1,1]$. This specific loss depends on the method used for distillation which can be either DERL, OUTL, or SOB. (c) the initial condition loss (c) The boundary condition loss for the new extended region $0,0.7]\times[-1,1]$. In this step, we consider two possibilities: we load the weights from the model of step 1, which we indicate with continual, or we start from a randomly initialized network, which we indicate with from-scratch. Finally, we consider a model that loads the weights of step $1$ but has no distillation term for the lower-left quadrant and should forget that region: we call this model PINN with no distillation. Either way, we consider the final model to be a predictor for the solution $u$ in the cumulative interval $0,0.7]\times[-1,1]$. We again save a dataset of its evaluation on the collocation points of the lower half of the domain.
@@ Results.
We now provide additional results for the transferring experiment on larger timespans in Section 5.1. Figure 20 shows the prediction error in the full time-space domain $(t,x)\in[0,1]\times[-1,1]$. We clearly see how DERL outperforms the model, especially compared to OUTL and the standard PINN, which converges early to a suboptimal solution.
E.4 Transfer to new domain
@@
The domain extension task we conducted shows how DERL can be leveraged to improve an existing physical model by extending its working domain. DERL allows for the transfer of physical information from a teacher model to a student one that is currently training to solve an equation in a new region of the domain. This way, the student model should be a good predictor of the solution of the equation in a larger domain than its teacher. The equation of interest is the Allen-Cahn equation (AC) in table 8. The architecture of the model is the same as in Section 4.1, that is we always work with MLPs with $4$ layers of $50$ neurons. In this case, no tuning is performed. We simulate a setting where the solution is to be found with no prior knowledge, as we are interested in the model that best works out-of-the-box, without any particular training procedure. The full domain is $1,1]\times[-1,1]$, which we divide in the lower-left quadrant $1,0]\times[-1,0]$, the lower-right quadrant $0,1]\times[-1,0]$ and the upper section $1,1]\times[0,1]$ as in figure 1. We first sampled a total of $1000$ random collocation points in the full domain. We also calculated the boundary conditions for each sub-region and the full one. Our experiment consists of three main steps and here we provide an in-depth description of them.
@@ Results.
As seen in Section 5.2, the PINN model with no distillation performs poorly and forgets the information from the lower half of the domain, which can be noticed from the error in figure 21. We plot the error comparison of the other methodologies on $u$ and their PDE residuals in figures 22 and 23.
@@ E.5 Transfer to new parameters
. The architecture of the model is the same as in Section 4.1, that is we always work with MLPs with $4$ layers of $50$ neurons. The models take as input $(x,y,\xi_{1},\xi_{2})$ and predict $\hat{u}(x,y;\xi_{1},\xi_{2})$. In this case, no tuning is performed. We simulate a setting where the solution is to be found with no prior knowledge, as we are interested in the model that best works out-of-the-box, without any particular procedure. The domain is $1,1]\times[-1,1]$ and we consider a total of $25$ couples of parameters $(\xi_{1},\xi_{2})$: $10$ are used in the first step, $10$ in the second step and the rest for final testing purposes. We sample a total of $1000$ random collocation points which will be used in every step, together with all couples of parameters. Our experiment consists of two main steps:
@@
    A second model is trained for $100$ epochs with the Adam optimizer and another $100$ with the BFGS optimizer with a loss that consists of (a) a PINN loss term to find the solution to the parametric equation for the next $10$ couples of $\xi_{1},\xi_{2}$, (b) a distillation loss on the dataset from step 1, that is on previous values of the parameters. This specific loss depends on the method used for distillation which can be either DERL, OUTL, or SOB. (c) The boundary condition loss. We again consider two possibilities: we load the weights from the model of step 1, which we indicate with continual, or we start from a randomly initialized network, which we indicate with from-scratch. Other possibilities are as in E.4. We finally test the models on the last $5$ couple of parameters.
@@ Results.
Appendix F On the Effect of Noise and Step Size on DERL
In this Section, we experiment with empirical derivatives with different values for the step size $h$ and with white noise applied to the true derivative values, to simulate the effects of real measurements on DERL. We aim to show the robustness of DERL under realistic conditions. However, step size is not an issue in our experimental setup: as shown in Section 4.1 and in Appendix D.4.3, one can apply a cubic interpolation on the available data to calculate empirical derivatives with any step size. Furthermore, step sizes and noise are never an issue in the transfer learning setup of Section 5, where partial derivatives are always calculated on the teacher network, without any need for approximation.
We experiment with the Allen-Cahn equation in Section 4.1 for the in-domain generalization, and with the damped Pendulum experiment in Section 4.2 for the generalization to new parameters. We experiment with the step sizes $h=0.001,0.01,0.05$ and with Gaussian noise with $\sigma=0.001,0.01,0.05$. We also report the results of SOB with the same setup for a direct comparison of the two methods.
@@ Appendix F On the Effect of Noise and Step Size on DERL
Results for the Allen-Cahn experiment are available in Table 15, while results for the damped pendulum experiment are available in Table 16. As we can see, DERL is robust to different noise values and with increasing step size. We notice that similar patterns observed in DERL are also present with SOB, showing that DERL extracts all the available information from partial derivatives, even without accessing information on $u$ directly.

### v3

@@ Appendix C Additional Results and Experiments for Section 4
C.1 Allen-Cahn equation
@@ C.2 Continuity equation
The original solution was calculated using finite volumes (Ferziger and Peric, 2001) on a grid with $\Delta x=\Delta y=0.01$ in the domain [1.5,1.5][-1.5,1.5] and time discretization of $\Delta t=0.001$ for $t\in[0,10]$. For the points used during training, we downsampled the full grid by a factor of $10$ for each spatiotemporal coordinate. In the testing phase, we test the model on the full grid, without downsampling. This way, we can test which model can best learn the solution on a finer grid, while being consistent with the equation. The empirical derivatives are calculated with $h=0.01$ on each component.
@@
C.3 Navier-Stokes: Kovasznay flow
@@ C.4 Pendulum
C.4.1 Differentiability with respect to Initial Conditions
@@ C.4.2 HNN and LNN setup
C.4.3 Results with empirical derivatives
@@ C.4.4 Conservative pendulum
C.5 Parametric Allen-Cahn Equation
C.5.1 Experiment with One parameter
Setup.
The MLPs have $4$ hidden layers with $50$ units each, similar to Raissi et al. (2019). The models take as input the coordinate and parameter of the equation $(x,y,\xi)$ and output the solution for the Allen-Cahn equation with parameter $\xi$, that is $\hat{u}(x,y;\xi)$. The networks are trained with batch size $100$ with the Adam optimizer for $50$ epochs, then are further optimized for another $50$ epochs with the BFGS optimizer on the whole dataset, similarly to Penwarden et al. (2023). The solution on $1000$ random collocation points is calculated for $16$ values of $\xi$. A train/validation/test split is applied, with $8$ parameters for training, $2$ for validation, and $6$ for testing. Testing values of $\xi$ are chosen to explore regions of the parameter space not seen during training and validation. The training and validation datasets are used for tuning the hyperparameters of the models, that is the learning rate and the weights of the loss terms in equation 17.
@@ Results.
C.5.2 Experiment with Two Parameters
@@
This Section provides additional results and information on the setup for the double parametric Allen-Cahn experiment in Section 4.2. The MLPs have $4$ hidden layers with $50$ units each, similar to Raissi et al. (2019). The models take as input the coordinate and parameters of the equation $(x,y,\xi_{1},\xi_{2})$ and output the estimate of $u$ with parameters $(\xi_{1},\xi_{2})$, that is $\hat{u}(x,y;\xi_{1},\xi_{2})$. The networks are trained with batch size $100$ with the Adam optimizer for $100$ epochs, then are further optimized for another $100$ epochs with the BFGS optimizer on the whole dataset, similarly to Penwarden et al. (2023). The solution on $1000$ random collocation points is calculated for $25$ couples of $(\xi_{1},\xi_{2})$. A train/validation/test split is applied, with $15$ couples of parameters for training, $5$ for validation, and $5$ for testing. The training and validation datasets are used for tuning the hyperparameters of the models, that is the learning rate and the weights of the loss terms in equation 17.
@@
Appendix D Additional Results and Experiments for Section 5
D.1 Higher Order Derivatives
@@
We consider the Korteweg-de Vries (KdV) equation, a third-order non-linear PDE (equation (KdV) in table 8), with $\nu=0.0025$. The IC is $u(0,x)=\cos(\pi x)$ with periodic BC for $u$ and $u_{x}$. The reference solution is first obtained using the Scipy (Virtanen et al., 2020) solver with Fast Fourier Transforms on a grid with $\Delta x=\Delta t=0.005$. Then, a PINN is carefully trained on equation (KdV) in table 8 and is treated as the teacher model to distill knowledge from. For the teacher model, we used an MLP with $9$ layers of $50$ units, $\tanh$ activation, trained for $200$ epochs with batch size $64$ and Adam optimizer. The higher number of layers is applied due to the complexity of the KdV task, as in Jagtap et al. (2020). After the teacher’s training, we save a dataset of its outputs $\hat{u}_{T}(x,t)$, its derivatives $\mathrm{D}\hat{u}_{T}(x,t)$, and its hessian matrix $\mathrm{D}^{2}\hat{u}_{T}(x,t)$ on the training collocation points. As always, partial derivatives are calculated via AD Baydin et al. (2018). Then, the student networks are trained with the specific distillation loss of the method, e.g. $\mathrm{D}\hat{u}_{S}-\mathrm{D}\hat{u}_{T}$ for DERL, and the original BC and IC. Since this is a third-order PDE, we are also interested in understanding the impact of higher-order derivatives on the performance, both in terms of the PDE residual and in terms of matching the true solution $u$ or the BC. We implemented Hessian learning (HESL), which learns the Hessian matrices $\mathrm{D}^{2}\hat{u}_{T}$ of the teacher model. We also explored its combination with DERL (DER+HESL, which learns both $\mathrm{D}\hat{u}_{T}$ and $\mathrm{D}^{2}\hat{u}_{T}(x,t)$) and Sobolev (SOB+HESL). In this last case, we use the approximation of the Hessian matrix as suggested in Czarnecki et al. (2017), while for HESL we use the full Hessian of the network, at the cost of an increase in computational time (see Section A.1 for further details). Each methodology has been tuned individually to find the hyperparameters that provide the best approximation of the true teacher’s solution $\hat{u}_{T}$.
@@ Results.
D.2 Neural Conservation Laws
@@
D.3 Transfer to a Larger Timespan
@@
    A second model is trained for $30.000$ steps and the Adam optimizer with the same loss as in equation 7 that consists of (a) a PINN loss term to find the solution to the equation in the second interval $0.35,0.7]\times[-1,1]$, (b) a distillation loss on the dataset from step 1, that is on points of the region $0,0.35]\times[-1,1]$. This specific loss depends on the method used for distillation, which can be either DERL, OUTL, or SOB. (c) The initial condition loss (c). The boundary condition loss for the new extended region $0,0.7]\times[-1,1]$. In this step, we consider two possibilities: we load the weights from the model of step 1, which we indicate with continual, or we start from a randomly initialized network, which we indicate with from-scratch. Finally, we consider a model that loads the weights of step $1$ but has no distillation term for the lower-left quadrant and should forget that region: we call this model PINN with no distillation. Either way, we consider the final model to be a predictor for the solution $u$ in the cumulative interval $0,0.7]\times[-1,1]$. We again save a dataset of its evaluation on the collocation points of the lower half of the domain.
@@
Figure 21: Prediction errors for the KdV transfer experiment. Vertical lines represent when a new step of incremental training is initiated.
@@ Results.
We now provide additional results for the transferring experiment on larger timespans in Section 5.1. Figure 20 shows the prediction error in the full time-space domain $(t,x)\in[0,1]\times[-1,1]$. We clearly see how DERL outperforms the model, especially compared to OUTL and the standard PINN, which converges early to a suboptimal solution. Finally, Figure 21 shows the prediction errors on the whole domain as a function of the training steps for all the methods. We see that the results of from-scratch are similar to the ones of the continual approach, with the latter performing slightly better.
D.4 Transfer to new domain
@@
The domain extension task we conducted shows how DERL can be leveraged to improve an existing physical model by extending its working domain. DERL allows for the transfer of physical information from a teacher model to a student one that is currently training to solve an equation in a new region of the domain. This way, the student model should be a good predictor of the solution of the equation in a larger domain than its teacher. The equation of interest is the Allen-Cahn equation (AC) in table 8. The architecture of the model is the same as in Section 4.1, that is we always work with MLPs with $4$ layers of $50$ neurons, similar to (Raissi et al., 2019). In this case, no tuning is performed. We simulate a setting where the solution is to be found with no prior knowledge, as we are interested in the model that best works out-of-the-box, without any particular training procedure. The full domain is $1,1]\times[-1,1]$, which we divide in the lower-left quadrant $1,0]\times[-1,0]$, the lower-right quadrant $0,1]\times[-1,0]$ and the upper section $1,1]\times[0,1]$ as in figure 1. We first sampled a total of $1000$ random collocation points in the full domain. We also calculated the boundary conditions for each sub-region and the full one. Our experiment consists of three main steps and here we provide an in-depth description of them.
@@ Results.
As seen in Section 5.2, the PINN model with no distillation performs poorly and forgets the information from the lower half of the domain, which can be noticed from the error in figure 22. We plot the error comparison of the other methodologies on $u$ and their PDE residuals in figures 23 and 24.
@@ D.5 Transfer to new parameters
. The architecture of the model is the same as in Section 4.1, that is we always work with MLPs with $4$ layers of $50$ neurons, similarly to Raissi et al. (2019). The models take as input $(x,y,\xi_{1},\xi_{2})$ and predict $\hat{u}(x,y;\xi_{1},\xi_{2})$. In this case, no tuning is performed. We simulate a setting where the solution is to be found with no prior knowledge, as we are interested in the model that best works out-of-the-box, without any particular procedure. The domain is $1,1]\times[-1,1]$, and we consider a total of $25$ couples of parameters $(\xi_{1},\xi_{2})$: $10$ are used in the first step, $10$ in the second step, and the rest for final testing purposes. We sample a total of $1000$ random collocation points which will be used in every step, together with all couples of parameters. Our experiment consists of two main steps:
@@
    A second model is trained for $100$ epochs with the Adam optimizer and another $100$ with the BFGS optimizer with a loss that consists of (a) a PINN loss term to find the solution to the parametric equation for the next $10$ couples of $\xi_{1},\xi_{2}$, (b) a distillation loss on the dataset from step 1, that is on previous values of the parameters. This specific loss depends on the method used for distillation, which can be either DERL, OUTL, or SOB. (c) The boundary condition loss. We again consider two possibilities: we load the weights from the model of step 1, which we indicate with continual, or we start from a randomly initialized network, which we indicate with from-scratch. Other possibilities are as in D.4. We finally test the models on the last $5$ couple of parameters.
@@ Results.
Appendix E On the Effect of Noise and Step Size on DERL
In this Section, we experiment with empirical derivatives with different values for the step size $h$ and with white noise applied to the true derivative values, to simulate the effects of real measurements on DERL. We aim to show the robustness of DERL under realistic conditions. However, step size is not an issue in our experimental setup: as shown in Section 4.1 and in Appendix C.4.3, one can apply a cubic interpolation on the available data to calculate empirical derivatives with any step size. Furthermore, step sizes and noise are never an issue in the transfer learning setup of Section 5, where partial derivatives are always calculated on the teacher network, without any need for approximation.
We experiment with the Allen-Cahn equation in Section 4.1 for the in-domain generalization, and with the damped Pendulum experiment in Section 4.2 for the generalization to new parameters. We experiment with the step sizes $h=0.001,0.01,0.05$ and with Gaussian noise with $\sigma=0.001,0.01,0.05$. We also report the results of SOB and OUTL+PINN (only for the noise, as empirical derivatives are not used) with the same setup for a direct comparison of the two methods.
@@ Appendix E On the Effect of Noise and Step Size on DERL
OUTL+PINN $h=0.05$ 0.05277 0.1848 0.1769 1.054 1.540
Results for the Allen-Cahn experiment are available in Table 15, while results for the damped pendulum experiment are available in Table 16. As we can see, DERL is robust to different noise values and with increasing step size. We notice that similar patterns observed in DERL are also present with SOB and OUTL+PINN, which is the least robust by far, showing that DERL extracts all the available information from partial derivatives, even without accessing information on $u$ directly.